OpenStudy (haisey_may):

Solve: 14x=7*10 (WILL MEDAL CORRECT ANSWER AND FAN)

OpenStudy (miss.rose):

X=5

OpenStudy (haisey_may):

Are you sure?

OpenStudy (haisey_may):

@Miss.Rose

OpenStudy (miss.rose):

yes

OpenStudy (haisey_may):

can you tell me why it is 5?

OpenStudy (miss.rose):

try it like this: 7*10= 70 14*5=70 so x=5

OpenStudy (haisey_may):

okay, but why would it be 70? (sorry btw. i really suck at math)

OpenStudy (miss.rose):

Well, as you know 7*10=70 so meaning what ever 14 is multiplied by it has to equal 70 in order for the problem to be true.

OpenStudy (haisey_may):

Oh okay. makes sense now! Thank you! @Miss.Rose

OpenStudy (miss.rose):

Your welcome.